how would i solve this problem? (z+5)^3
This is what I have so far (z+5)(z+5)(z+5)
I am lost.

Answer by lwsshak3(11628)

how would i solve this problem? (z+5)^3
(z+5)^3=(z+5)(z+5)(z+5)=(z+5)(z+5)^2=(z+5)(z^2+10z+25)
multiply by z, then multiply by 5, then combine like terms
=z^3+10z^2+25z+5z^2+50z+125
=z^3+15z^2+75z+125